OpenStudy (anonymous):

@amistre64 @Hero @ganeshie8 @phi @.Sam. @mathstudent55 HELP MEDALS Use signed numbers to solve. A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. What was the net gain or loss in altitude? A. lost 2450 ft B. lost 850 ft C. gained 850 ft D. gained 2450 ft

OpenStudy (anonymous):

@Reaper534

OpenStudy (gir_lover_♥):

Its subtracted then added then subratcted

OpenStudy (anonymous):

thanks

OpenStudy (gir_lover_♥):

Np

OpenStudy (anonymous):

wait so it is 1200-800+450

OpenStudy (amistre64):

start at 0 ....

OpenStudy (anonymous):

am i right or wrong

OpenStudy (anonymous):

-1200 + 800=-400-450=-850

OpenStudy (anonymous):

from -1200 to -850 did it go higher or lower?

OpenStudy (anonymous):

lower

OpenStudy (anonymous):

lower

OpenStudy (anonymous):

B. right?

OpenStudy (anonymous):

@Reaper534 am i right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you went higher

OpenStudy (anonymous):



OpenStudy (anonymous):

wait so what is the answer then i am confused

OpenStudy (anonymous):

are you sure those are the correct answer choices?

OpenStudy (anonymous):

I think its C then @amistre64 what do you think?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with more